# Magic tricycles: efficient magic state generation with finite block-length quantum LDPC codes

### v1

@@ 2.5 Implementation of syndrome extraction circuits
We construct optimal-depth syndrome extraction circuits for the tricycle codes presented. Denote the data qubit sectors as $D_{1}$, $D_{2}$, $D_{3}$, the Z-check sectors as $Z_{1}$, $Z_{2}$, $Z_{3}$, and the X-check sector as $X$. Our construction is displayed in Fig.˜5a, where each CNOT symbol represents a parallel group of physical CNOTs between a check sector and a data sector, associated to a specific permutation. For example, the first CNOT symbol corresponds to CNOTs from the $i$-th X-check to the $j$-th qubit in $D_{3}$ whenever the $(i,j)$ entry of $\mathbf{C}_{1}^{T}$ is 1. Since $\mathbf{C}_{1}^{T}$ is a permutation matrix, there is one and only one $j$ for each $i$. The circuit achieves the optimal CNOT depth of 12. For a proof of correctness and applications of this construction in other settings, see Appendix˜F.
@@
Entangling gates are performed by exciting adjacent pairs of qubits to Rydberg states. Parallel CNOTs between corresponding qubits in two sectors are realized by overlaying those sectors and applying a global Rydberg laser pulse, as shown in Fig.˜5c (dots and squares denote data and check qubits, respectively. As the colors suggests, the check sector has been permuted with $x$-cycling by one unit and $y$-cycling by two units.) To prevent unintended interactions between qubits within the same sector, the pairs are separated by a sufficient distances determined by atom species and Rydberg state.
@@
Executing one syndrome cycle requires a constant number of AOD movements. Let $t_{\text{wsp}}$ and $t_{\text{ent}}$ denote the time for AOD to traverse a sector in the workspace and the entangling zone, respectively. In the implementation above, permutation, fetch, and put-back for all $Z$ sectors can be done in parallel. Each CNOT layer then involves two fetch and two put-back operations, each bounded by $3t_{\text{ent}}$ (say, from $X$ to $D_{1}$). Permutations for $Z$ and $X$ sectors may require $x$-, $y$-, and $z$-cycling, each up to $2t_{\text{wsp}}$. Thus, the per-layer duration is bounded by $12(t_{\text{wsp}}+t_{\text{ent}})$. Further details and possible optimization are discussed in Appendix˜G.
@@ 3 Discussion and Outlook
We have introduced a class of quantum LDPC codes—tricycle codes—that support constant-depth, single-shot preparation of logical magic states and show strong resilience to realistic circuit-level noise. We showed that these codes allow compact, constant depth preparation of high-magic states known as hypergraph magic states while maintaining competitive circuit-level noise suppression and a high threshold of $0.4\%$. These hypermagic states in turn embed individually distillable $\overline{CCZ}$-type magic states, enabling universal quantum computation by using tricycle codes as magic state factories. The promising performance of these codes can be improved further along several directions. The asymmetry between the $X$ and $Z$ sectors of tricycle codes make them particularly advantageous for leveraging noise bias in entangling gate operations—a feature pervasive in leading experimental platforms [1, 5, 64, 79, 11]. Additionally, loss and leakage errors—prevalent in many hardware architectures—represent another avenue for enhancing performance; recent work has shown that appropriate handling of such errors can be leveraged to further improve logical error rates [2, 5]. While our present analysis focuses on unbiased depolarizing noise and neglects loss errors, future studies integrating tailored noise bias exploitation and explicit loss-/leakage-leveraging strategies with these codes are likely to yield further gains and represent a compelling direction for subsequent research.
@@
Beyond their standalone utility as distillation factories, integrating tricycle codes into broader quantum architectures presents several important challenges. A key open problem is the seamless injection or teleportation of distilled magic states into computational code blocks, such as those based on high-performance bicycle codes. While lattice surgery offers a robust, code-agnostic method for logical state transfer—and has seen significant theoretical and experimental development [25, 95, 93, 45, 85, 27, 26]—a particularly promising alternative is transversal teleportation based on natural isomorphisms between tricycle and bicycle codes. Similar to teleportation protocols between 3D and 2D color codes [84, 28], such schemes using transversal one-way CNOT gates between $3D$ and $2D$ product codes [42] may enable direct, fault-tolerant transfer of magic states. In particular, we believe this should be possible between bivariate-bicycle codes and bivariate-tricycle codes or between triviariate-bicycle codes and trivariate-tricycle codes, as they share the same product structure. Using extensions of the analytical tools in Ref. [30], we believe it should be possible to show that the logical operators of pairs of some such $2D$ and $3D$ group-algebra codes split into isomorphic sectors as in hypergraph product codes [90] – a natural setting for efficient gate teleportation with transversal one-way CNOTs. If realized, these protocols could further reduce space-time overhead and enable highly modular architectures with efficiently integrated magic state distillation.
@@
While we employed the cup-product-gate based construction of Ref. [16] to construct transversal $CCZ$ circuits, alternative approaches may yield shorter-depth circuits or simpler logical structure. Another important direction is improving the yield of disjoint $\overline{CCZ}$ gates that can be extracted from the hypergraph magic states produced by these circuits (Appendix˜C). This corresponds to the problem of finding the subrank of a binary tensor [24, 48, 35, 55], which is computationally hard. While we used mixed-integer programming, new heuristic strategies could uncover larger values of $K_{CCZ}$ for the codes we consider. In parallel, a deeper understanding of the structure of these high-magic hypergraph states—produced by the transversal $CCZ$ circuits described in Sec.˜2.2—may allow one to compile them directly into useful quantum circuits, offering an alternative to extracting only disjoint $\overline{CCZ}$ gates.
@@ 4 Acknowledgments
We thank Nazli Ugur Koyluoglu, Andi Gu, Rohith Sajith, Nishad Maskara, Shayan Majidy, Hengyun Zhou, Harry Putterman, Qian Xu, Brenden Roberts, Jin Ming Koh, Andrei Diaconu, and Prof. Jason Cong for valuable discussions. We acknowledge financial support from IARPA and the Army Research Office, under the Entangled Logical Qubits program (Cooperative Agreement Number W911NF-23-2-0219), the DARPA MeasQuIT program (grant number HR0011-24-9-0359), the Center for Ultracold Atoms (a NSF Physics Frontiers Center, PHY-1734011), the National Science Foundation (grant numbers PHY-2012023 and CCF-2313084), the Wellcome Leap Quantum for Bio program, Harvard Quantum Initiative Postdoctoral Fellowship (DBT) and QuEra Computing. After the completion of this project, we became aware of related work studying tricycle codes, albeit with less focus on transversal magic gates [47].
@@ Appendix A Code construction and properties
We focus on codes with $G$ equal to a product of one, two, or three cyclic groups ($k=1,2,3$). By analogy to the nomenclature for bicycle codes [49, 13, 65], these may be referred to as generalized tricycle, bivariate-tricycle, or trivariate-tricycle codes. We denote the generators by monomials $x_{1}\equiv x$, $x_{2}\equiv y$, and $x_{3}\equiv z$. The polynomials and groups used in the examples of footnote˜2 are listed in LABEL:table::code_polys.
Table 4: Polynomials used to construct codes in footnote˜2 corresponding to elements of the group algebra of $G=\mathbb{Z}_{l}\times\mathbb{Z}_{m}\times\mathbb{Z}_{n}$. If there are fewer than three cyclic group factors, the corresponding element in the tuple of group orders $(l,m,n)$ is left empty.
@@
Consider a connected three-block Abelian group algebra code constructed from group algebra elements $a,b,c\in\mathbb{F}_{2}[G]$, with corresponding binary matrices $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{F}_{2}^{n\times n}$ where $n=|G|$. Further suppose that the classical binary linear codes with parity check matrices $\mathbf{A^{T}},\mathbf{B^{T}},\mathbf{C^{T}}$ have minimum distances $d_{A}^{T},d_{B}^{T},d_{C}^{T}$. We use the convention that $d=\infty$ if the corresponding matrix is full rank. Let $N=G_{a}\cap G_{b}\cap G_{c}$ be the intersection subgroup as defined in Definition˜2. The minimum $Z$ distance of the quantum code is bounded below as
$d_{Z}\geq\frac{1}{|N|}min(d_{A}^{T},d_{B}^{T},d_{C}^{T})$ (15)
@@
This follows from the proof of Statement 10 in [54]. In particular, Lemma 3 of [54] also holds for three-dimensional homological products by decomposing them into a direct sum of three-dimensional homological products over cyclic group algebras (analogous to Appendix B. in [66]). The matrices $\mathbf{A},\mathbf{B},\mathbf{C}$ are constructed identically, and following the proof of Lemma 3. in [54], each component code over a cyclic group algebra can be shown to have all-unit Smith Normal Forms under the assumptions of the lemma. The rest of the proof is identical, applied instead to the matrix $H_{X}$ defined in Eq.˜1. ∎
@@
In the following conjecture, we argue that $d_{Z}\leq d_{X}$ for the tricycle codes. This would make Theorem˜A.1 a lower bound on both the $X$ and $Z$. Although we are unable to construct a formal proof, extensive numerical searches over several millions of random codes indeed validate this claim.
Conjecture 1.
All three-block Abelian group algebra codes defined by Eq.˜1 satisfy $d_{Z}\leq d_{X}$.
@@
[…]
@@
The quantum CSS code associated with the tricycle construction is defined by the first three terms of Eq.˜18, with consecutive spaces representing $X$-checks, qubits, and $Z$-checks respectively. Binarizing the coboundary maps yields the parity check matrices in Eq.˜1, and the final map defines the meta-checks in Eq.˜3. This complex appears in the appendix of [30], and tricycle codes of this form are also briefly noted in [16].

### v2

@@ 2.5 Implementation of syndrome extraction circuits
We construct optimal-depth syndrome extraction circuits for the tricycle codes presented. Denote the data qubit sectors as $D_{1}$, $D_{2}$, $D_{3}$, the Z-check sectors as $Z_{1}$, $Z_{2}$, $Z_{3}$, and the X-check sector as $X$. Our construction for 4444-4-4 codes is displayed in Fig.˜5a, where each CNOT symbol represents a parallel group of physical CNOTs between a check sector and a data sector, associated with a specific permutation. For example, the first CNOT symbol corresponds to CNOTs from the $i$-th X-check to the $j$-th qubit in $D_{3}$ whenever the $(i,j)$ entry of $\mathbf{C}_{1}^{T}$ is 1. Since $\mathbf{C}_{1}^{T}$ is a permutation matrix, there is one and only one $j$ for each $i$. The circuit achieves the optimal CNOT depth of 12. The construction consists of five stages represented by colors. For the 4424-4-2 and 4224-2-2 codes, the overall stage structure is preserved, although some stages contain fewer layers. The design allows certain degrees of freedom: specifically, how to assign the matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ to the stages, as well as the order of layers within each stage. These degrees of freedom lead to 1728 possible circuits for the 4224-2-2 codes. An exhaustive search across these 1728 circuits was performed, from which we select the circuit with the highest noise-suppression for each code in the circuit-noise simulation results of Sec.˜2.4. For the general construction and a proof of its correctness, see Appendix˜H.
@@
Entangling gates are performed by exciting adjacent pairs of qubits to Rydberg states. Parallel CNOTs between corresponding qubits in two sectors are realized by overlaying those sectors and applying a global Rydberg laser illumination, as shown in Fig.˜5c (dots and squares denote data and check qubits, respectively. As the colors suggests, the check sector has been permuted with $x$-cycling by one unit and $y$-cycling by two units.) To prevent unintended interactions between qubits within the same sector, the pairs are separated by a sufficient distances determined by atom species and Rydberg state.
@@
Executing one syndrome cycle requires a constant number of AOD movements. Let $t_{\text{wsp}}$ and $t_{\text{ent}}$ denote the time for AOD to traverse a sector in the workspace and the entangling zone, respectively. In the implementation above, permutation, fetch, and put-back for all $Z$ sectors can be done in parallel. Each CNOT layer then involves two fetch and two put-back operations, each bounded by $3t_{\text{ent}}$ (say, from $X$ to $D_{1}$). Permutations for $Z$ and $X$ sectors may require $x$-, $y$-, and $z$-cycling, each up to $2t_{\text{wsp}}$. Thus, the per-layer duration is bounded by $12(t_{\text{wsp}}+t_{\text{ent}})$. Further details and possible optimization are discussed in Appendix˜I.
@@ 3 Discussion and Outlook
We have introduced a class of quantum LDPC codes—tricycle codes—that support constant-depth, single-shot preparation of logical magic states and show strong resilience to realistic circuit-level noise with a high threshold of $0.5\%$. We showed that certain families of tricycle codes allow compact, constant depth preparation of high-magic states known as hypergraph magic states. These hypermagic states in turn embed individually distillable $\overline{CCZ}$-type magic states, enabling universal quantum computation by using tricycle codes as magic state factories. The promising performance of these codes can be improved further along several directions. The asymmetry between the $X$ and $Z$ sectors of tricycle codes make them particularly advantageous for leveraging noise bias in entangling gate operations—a feature pervasive in leading experimental platforms [1, 7, 74, 90, 13]. Additionally, loss and leakage errors—prevalent in many hardware architectures—represent another avenue for enhancing performance; recent work has shown that appropriate handling of such errors can be leveraged to further improve logical error rates [3, 7]. While our present analysis focuses on unbiased depolarizing noise and neglects loss errors, future studies integrating tailored noise bias exploitation and explicit loss-/leakage-leveraging strategies with these codes are likely to yield further gains and represent a compelling direction for subsequent research. Moreover, improved decoding methods would be beneficial as we observe over an order-of-magnitude gap in logical error rates between standard BP+OSD and BP+LSD decoders and an exact but exponentially expensive most-likely-error (MLE) decoder. Encouragingly, recent work has proposed new decoders that enhance performance while maintaining practical inference times [72, 12, 67, 68].
@@
Beyond their standalone utility as distillation factories, integrating tricycle codes into broader quantum architectures presents several important challenges. A key open problem is the seamless injection or teleportation of distilled magic states into computational code blocks, such as those based on high-performance bicycle codes. While lattice surgery offers a robust, code-agnostic method for logical state transfer—and has seen significant theoretical and experimental development [27, 107, 105, 49, 96, 30, 29, 4, 112, 28]—a particularly promising alternative is transversal teleportation based on natural isomorphisms between tricycle and bicycle codes. Similar to teleportation protocols between 3D and 2D color codes [95, 31], such schemes using transversal one-way CNOT gates between $3D$ and $2D$ product codes [45] may enable direct, fault-tolerant transfer of magic states. In particular, we believe this should be possible between bivariate-bicycle codes and bivariate-tricycle codes or between triviariate-bicycle codes and trivariate-tricycle codes, as they share the same product structure. Using extensions of the analytical tools in Ref. [33], we believe it should be possible to show that the logical operators of pairs of some such $2D$ and $3D$ group-algebra codes split into isomorphic sectors as in hypergraph product codes [102] – a natural setting for efficient gate teleportation with transversal one-way CNOTs. A version of this idea has recently been explored in [60]. If realized, these protocols could further reduce space-time overhead and enable highly modular architectures with efficiently integrated magic state distillation. We will explore a specific magic-state factory architecture based on high-rate tricycle codes with transversal teleportation onto a target high-rate bicycle code in a follow-up work.
@@
Another important direction is improving the yield of disjoint $\overline{CCZ}$ gates that can be extracted from the hypergraph magic states produced by these circuits (Appendix˜F). This corresponds to the problem of finding the subrank of a binary tensor [26, 52, 38, 63], which is computationally hard. While we used mixed-integer programming, new heuristic strategies could uncover larger values of $K_{CCZ}$ for the codes we consider. In parallel, a deeper understanding of the structure of these high-magic hypergraph states—produced by the transversal $CCZ$ circuits described in Sec.˜2.2—may allow one to compile them directly into useful quantum circuits, offering an alternative to extracting only disjoint $\overline{CCZ}$ gates.
@@ 4 Acknowledgments
We thank Qian Xu, Nazli Ugur Koyluoglu, Rohith Sajith, Nishad Maskara, Shayan Majidy, Hengyun Zhou, Harry Putterman, Brenden Roberts, Jin Ming Koh, Andrei Diaconu, and Jason Cong for valuable discussions. We particularly thank Qian Xu for detailed feedback on our results and manuscript. We acknowledge financial support from IARPA and the Army Research Office, under the Entangled Logical Qubits program (Cooperative Agreement Number W911NF-23-2-0219), the DARPA MeasQuIT program (grant number HR0011-24-9-0359), the Center for Ultracold Atoms (a NSF Physics Frontiers Center, PHY-1734011), the National Science Foundation (grant numbers PHY-2012023 and CCF-2313084), the Wellcome Leap Quantum for Bio program, Harvard Quantum Initiative Postdoctoral Fellowship (DBT) and QuEra Computing. After the completion of this project, we became aware of related work studying tricycle codes, albeit with less focus on transversal magic gates [51]. Another related work that appeared around the same time also introduced a general theory for multi-cyclic Abelian group-algebra codes [61].
@@ Appendix A Code construction and properties
We focus on codes with $G$ equal to a product of three cyclic groups ($k=3$) as we found these codes to generally have better parameters than those with $k=1,2$ or higher. By analogy to the nomenclature for bicycle codes [53, 15, 76], these may be referred to as trivariate-tricycle codes. The $k=1,2$ cases may be referred to as generalized tricycle and bivariate-tricycle respectively. We denote the generators by monomials $x_{1}\equiv x$, $x_{2}\equiv y$, and $x_{3}\equiv z$. The polynomials and groups used in the examples of footnote˜2 are listed in LABEL:table::code_polys.
Table 2: Polynomials used to construct trivariate-tricycle codes in footnote˜2 corresponding to elements of the group algebra of $G=\mathbb{Z}_{l}\times\mathbb{Z}_{m}\times\mathbb{Z}_{n}$. If there are fewer than three cyclic group factors, the corresponding element in the tuple of group orders $(l,m,n)$ is left empty.
@@
Consider a connected three-block Abelian group algebra code constructed from group algebra elements $a,b,c\in\mathbb{F}_{2}[G]$, with corresponding binary matrices $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{F}_{2}^{n\times n}$ where $n=|G|$. Further suppose that the classical binary linear codes with parity check matrices $\mathbf{A^{T}},\mathbf{B^{T}},\mathbf{C^{T}}$ have minimum distances $d_{A}^{T},d_{B}^{T},d_{C}^{T}$. We use the convention that $d=\infty$ if the corresponding matrix is full rank. Let $N=G_{a}\cap G_{b}\cap G_{c}$ be the intersection subgroup as defined in Definition˜2. The minimum distance of the quantum code is bounded below as
$d\geq\frac{1}{|N|}\min\{d_{A}^{T},d_{B}^{T},d_{C}^{T}\}$ (15)
@@
The proof that $d_{Z}$ satisfies this inequality follows from the proof of Statement 10 in [62]. In particular, Lemma 3 of [62] also holds for three-dimensional homological products by decomposing them into a direct sum of three-dimensional homological products over cyclic group algebras (analogous to Appendix B. in [77]). The matrices $\mathbf{A},\mathbf{B},\mathbf{C}$ are constructed identically, and following the proof of Lemma 3. in [62], each component code over a cyclic group algebra can be shown to have all-unit Smith Normal Forms under the assumptions of the lemma. The rest of the proof is identical, applied instead to the matrix $H_{X}$ defined in Eq.˜1.
Then, in Theorem˜A.2, we prove that $d_{Z}\leq d_{X}$, making Eq.˜15 a lower bound for the overall code distance. ∎
@@
In the following theorem, we show that tricycle codes satisfy $d_{Z}\leq d_{X}$s. This would make Theorem˜A.1 a lower bound on both the $X$ and $Z$ distances.
Theorem A.2.
All connected three-block Abelian group algebra codes defined by Eq.˜1 satisfy $d_{Z}\leq d_{X}$.
Proof.
We will assume that none of $A_{1}$, $B_{1}$, $C_{1}$ are full rank, else the tricycle code is trivial with $k=0$ and $d=\infty$ by convention. Noting $R\cong\bigoplus_{g\in N}\mathbb{F}_{2}$, the map $\mathbbm{b}_{N}:R\rightarrow\mathbb{F}_{2}{|N|}$ defined by $\mathbbm{b}_{N}:a=\sum_{g}a_{g}g\mapsto(a_{g_{1}}\ \cdots\ a_{g_{|}N|})$ is a vector space isomorphism. With action defined element-wise, it can be extended to $\mathbbm{b}_{N}^{m}:R^{m}\rightarrow\mathbb{F}_{2}^{m}$; we omit the superscript hereafter. We define the norm $|\cdot|_{R}$ on $\mathbb{R}^{m}$ via $|v|_{R}=\sum_{i=1}^{m}|\{g\in G:(v_{i})_{g}=1\}|$ for any $v\in R^{m}$. Importantly, $|v|_{R}=|\mathbbm{b}_{N}(v)|$ in general. Let $\mathbb{B}_{N}:R\rightarrow\mathbb{F}_{2}^{|N|\times|N|}$ be the regular representation of $N$ introduced in Eq.˜4, extended element-wise and linearly to $R^{m\times m}$. For any $A\in R^{m\times m}$ and $v\in R^{m}$, $\mathbbm{b}_{N}(Av)=\mathbb{B}_{N}(A)\mathbbm{b}_{N}(v)$.
Let $x=(x_{1}\ x_{2}\ x_{3})\in R^{l_{a}l_{b}l_{c}}$ denote a minimum weight logical $X$ operator. Define
$(\delta^{-1})^{T}:=\begin{pmatrix}A_{1}^{T}\otimes_{R}I_{l_{b}}&I_{l_{a}}\otimes_{R}B_{1}^{T}\end{pmatrix}\hskip 20.0pt\delta^{0}:=\begin{pmatrix}I_{l_{a}}\otimes_{R}B_{1}&A_{1}\otimes_{R}I_{l_{b}}\end{pmatrix}$. (16)
[…]
Expressing $H_{X}/H_{Z}$ in terms of $(\delta^{-1})^{T}/\delta^{0}$,
$H_{X}=\begin{pmatrix}(\delta^{-1})^{T}\otimes_{R}I_{l_{c}}&I_{l_{a}l_{b}}\otimes_{R}C_{1}^{T}\end{pmatrix}\hskip 20.0ptH_{Z}=\begin{pmatrix}\delta^{0}\otimes_{R}I_{l_{c}}&0\\ I_{2l_{a}l_{b}}\otimes_{R}C_{1}&\delta^{-1}\otimes_{R}I_{l_{c}}\end{pmatrix}$. (17)
From the Künneth theorem for $R$-module cohomology (see section Homological products of the supplementary material and [65]),
$\frac{\ker H_{X}}{\operatorname{\text{im}}H_{Z}^{T}}\cong\left(\frac{R^{l_{a}}}{\operatorname{\text{im}}(\delta^{-1})^{T}}\otimes_{R}\ker C_{1}^{T}\right)\oplus\left(\frac{\ker(\delta^{-1})^{T}}{\operatorname{\text{im}}\delta^{0T}}\otimes_{R}\frac{R^{l_{c}}}{\operatorname{\text{im}}C_{1}^{T}}\right).$ (18)
[…]
Repeating the analysis for the other two ways to express the tricycle code as a $R$-homological product of a bicycle code and a classical group-algebra code, the above analysis is only incomplete in the case
$x\in\operatorname{\text{im}}\begin{pmatrix}T_{A}&0\\ 0&I_{l_{a}l_{b}l_{c}}\\ T_{C}&0\end{pmatrix}\cap\operatorname{\text{im}}\begin{pmatrix}T_{A}&0\\ T_{B}&0\\ 0&I_{l_{a}l_{b}l_{c}}\end{pmatrix}\cap\operatorname{\text{im}}\begin{pmatrix}0&I_{l_{a}l_{b}l_{c}}\\ T_{B}&0\\ T_{C}&0\end{pmatrix}$, (19)
where $T_{A}=A_{1}\otimes_{R}I_{l_{b}}\otimes_{R}I_{l_{c}}$, $T_{B}=I_{l_{a}}\otimes_{R}B_{1}\otimes_{R}I_{l_{c}}$, and $T_{C}=I_{l_{a}}\otimes_{R}I_{l_{b}}\otimes_{R}C_{1}$. Such an $x$ exists iff there exist $u_{1},u_{2},u_{3}\in R^{2l_{a}l_{b}l_{c}}$ such that $x_{1}=T_{A}u_{1}=T_{A}u_{2}$, $x_{2}=T_{B}u_{2}=T_{B}u_{3}$, and $x_{3}=T_{C}u_{1}=T_{C}u_{3}$, yet
$x\notin\operatorname{\text{im}}H_{X}^{T}=\operatorname{\text{im}}\begin{pmatrix}T_{A}\\ T_{B}\\ T_{C}\end{pmatrix}$. (20)
Let $v_{A}=u_{1}+u_{2}$, $v_{B}=u_{2}+u_{3}$, so $u_{1}+u_{3}=v_{A}+v_{B}$. It follows that $v_{A}\in\ker T_{A},v_{B}\in\ker T_{B},v_{A}+v_{B}\in\ker T_{C}$. Expressing
$x=\begin{pmatrix}T_{A}u_{1}\\ T_{B}u_{2}\\ T_{C}u_{3}\end{pmatrix}=\begin{pmatrix}T_{A}u_{1}\\ T_{B}(u_{1}+v_{A})\\ T_{C}(u_{1}+v_{A}+v_{B})\end{pmatrix}=\begin{pmatrix}T_{A}u_{1}\\ T_{B}(u_{1}+v_{A})\\ T_{C}u_{1}\end{pmatrix}=H_{X}^{T}u_{1}+\begin{pmatrix}0\\ T_{B}v_{A}\\ 0\end{pmatrix}$, (21)
a solution for $x$ exists iff $(0\ T_{B}v_{A}\ 0)$ is a nontrivial logical $X$ operator, which requires $v_{A}\notin\ker T_{B}$. Note $v_{A}\in\ker T_{A}\cap(\ker T_{B}+\ker T_{C})$. Now we use the following lemma.
Lemma A.3.
In this setting, the intersection can be distributed over subspace addition.
$\ker T_{A}\cap(\ker T_{B}+\ker T_{C})=(\ker T_{A}\cap\ker T_{B})+(\ker T_{A}\cap\ker T_{C})$ (22)
Proof.
The reverse inclusion is generally true. For the forward direction, define the canonical projection $\Pi:R^{l_{a}l_{b}l_{c}}\rightarrow\ker T_{A}$. Any $v\in\ker T_{A}\cap(\ker T_{B}+\ker T_{C})$ satisfies $\Pi v=v$ and can be written as $v=t_{B}+t_{C}$, where $t_{B}\in\ker T_{B}$ and $t_{c}\in\ker T_{C}$. Since $\Pi$ factorizes as $\Pi=\pi\otimes_{R}I_{l_{b}}\otimes_{R}I_{l_{c}}$, where $\pi:R^{l_{a}}\rightarrow\ker A_{1}$, $\Pi t_{B}\in\ker T_{B}$ and $\Pi t_{C}\in\ker T_{C}$. By definition, $\Pi t_{B},\Pi t_{C}\in\ker T_{A}$. ∎
Thus, $v_{A}\in(\ker T_{A}\cap\ker T_{B})+(\ker T_{A}\cap\ker T_{C})\subseteq\ker T_{B}+(\ker T_{A}\cap\ker T_{C})$. If we write $v_{A}\in t_{B}+t_{AC}$, where $t_{B}\in\ker T_{B}$, $t_{AC}\in\ker T_{A}\cap\ker T_{C}$, then, observe that $(0\ T_{B}v_{A}\ 0)=H_{X}^{T}t_{AC}$ and is thus logically trivial. This contradicts Eq.˜20, so Eq.˜19 cannot hold. ∎
Corollary A.3.1.
Theorem˜A.2 extends to disconnected codes.
Proof.
Any disconnected tricycle code is a direct sum of connected tricycle codes. Theorem˜A.2 can then be applied to the component with the lowest weight logical $X$ operator. ∎
@@
$\mathbb{F}_{2}[G]$$\mathbb{F}_{2}[G]^{3}$$\mathbb{F}_{2}[G]^{3}$$\mathbb{F}_{2}[G]$$\scriptstyle{\left(\begin{smallmatrix}a\\ b\\ c\end{smallmatrix}\right)}$$\scriptstyle{\left(\begin{smallmatrix}c&0&a\\ 0&c&b\\ b&a&0\end{smallmatrix}\right)}$$\scriptstyle{\left(\begin{smallmatrix}b&a&c\end{smallmatrix}\right)}$ (25)
@@
The quantum CSS code associated with the tricycle construction is defined by the first three terms of Eq.˜25, with consecutive spaces representing $X$-checks, qubits, and $Z$-checks respectively. Binarizing the coboundary maps yields the parity check matrices in Eq.˜1, and the final map defines the metachecks in Eq.˜3. This complex appears in the appendix of [33], and tricycle codes of this form are also briefly noted in [18].